If A and B are independent events with P(A) = 0.7 and P(B) = 0.8, then P(A∪B) = .................... .
0.94
Use the addition rule: P(A∪B) = P(A) + P(B) - P(A∩B).
For independent events P(A∩B) = P(A) x P(B) = 0.7 x 0.8 = 0.56.
So P(A∪B) = 0.7 + 0.8 - 0.56 = 1.5 - 0.56 = 0.94.
Alternatively, P(A∪B) = 1 - P(neither) = 1 - (0.3 x 0.2) = 1 - 0.06 = 0.94. Hence, the answer is 0.94.
The score in a certain examination follows normal distribution with mean 50 and standard deviation 16 then the marks of top 1% candidates will have score more than .................. .
Suppose the number of accidents per week on a certain highway follows Poisson distribution with mean 2 then the percentage of weeks having no accidents is ....................
(Given e-2 = 0.1353)
A box contains 8 white and 3 red balls. If two balls are drawn at random, find the probability that one is white and one is red.
For a normal curve with mean of 120 and standard deviation of 35, what proportion (in percent) of the area under the curve will lie between the values of 85 and 155 ?
A book contains 600 typographic errors in its 600 pages. These typo errors occur completely at random. Using the Poisson distribution, the probability, that a page contains no error, is :